# On-device Anomaly Detection in Conveyor Belt Operations

### v2

@@ V Conclusion
This work contributes to Mining 4.0 by proposing two novel approaches for real-time anomaly detection in the continuous operation of hydraulic conveyor belt systems using MCUs. These pattern recognition methods combine manual feature extraction, threshold-based cycle detection, and supervised TinyML classifiers to accurately identify normal and abnormal duty cycles. By evaluating the performance of different approaches and ML classifiers across two distinct datasets, we identified the most effective methods for achieving accurate and generalizable classification. These methods are essential for conveyor belt system monitoring and predictive maintenance. The first approach performed better when trained and tested on the same dataset, achieving an F1-score of 97.3% for normal and 80.2% for abnormal cycles. In contrast, the second approach showed superior performance on an independent dataset, with F1-scores of 91.3% for normal and 67.9% for abnormal cycles. This comparison highlights both the strengths and adaptability challenges of the methods across varying operational contexts. Furthermore, both approaches outperformed the state-of-the-art method on both datasets, validating their suitability for real-world scenarios.
When deployed in two resource-constrained MCUs, the methods demonstrated stable performance with 32-bit float representation and even slight improvements with 8-bit integer quantization, resulting from minimal quantization error. Energy consumption remained low, at 13.3 $\mu$J and 20.6 $\mu$J per inference on a Raspberry Pi Pico, enabling seamless integration into conveyor control systems and potential for automated preventive actions.

### v3

@@ V Conclusion
This work contributes to Mining 4.0 by proposing two novel approaches for real-time anomaly detection in the continuous operation of hydraulic conveyor belt systems using MCUs. These pattern recognition methods combine manual feature extraction, threshold-based cycle detection, and supervised TinyML classifiers to accurately identify normal and abnormal duty cycles. By evaluating the performance of different approaches and ML classifiers across two distinct datasets, we identified the most effective methods for achieving accurate and generalizable classification. These methods are essential for conveyor belt system monitoring and predictive maintenance. The first alternative (Approach-2) performed better when trained and tested on the same dataset, achieving an F1-score of 97.3% for normal and 80.2% for abnormal cycles. In contrast, the second alternative (Approach-3) showed superior performance on an independent dataset, with F1-scores of 91.3% for normal and 67.9% for abnormal cycles. This comparison highlights both the strengths and adaptability challenges of the methods across varying operational contexts. Furthermore, both approaches outperformed the state-of-the-art method on both datasets, validating their suitability for real-world scenarios.
When deployed in two resource-constrained MCUs, the methods demonstrated stable performance with 32-bit float representation and even slight improvements with 8-bit integer quantization, resulting from minimal quantization error. Energy consumption remained low, at 13.3 µJ and 20.6 µJ per inference on a Raspberry Pi Pico, enabling seamless integration into conveyor control systems and potential for automated preventive actions.